OpenStudy (anonymous):

I need some serious help! Solve S=4v^2 for v Solve M=2x+3y for y Solve P= 2(l+w) for l

OpenStudy (chaise):

S=4v^2 for v S/4 = v^2 v = \[\sqrt{s/4}\] M = 2x+3y for y m - 2x = 3y (m-2x)/3 = y Solve P= 2(l+w) for l P = 2l + 2w p - 2w = 2l (p-2w)/2 = L Need any further help with how these answers where obtained?

OpenStudy (anonymous):

I believe you can simplify the first one to just sqrtS/2

OpenStudy (anonymous):

with a +/_ in front :)

OpenStudy (anonymous):

yes you cannot forget that, like i just did. :)

OpenStudy (anonymous):

oh wow guys! thanks! i think that's it! You guys helped a lot! :D

OpenStudy (chaise):

my pleasure.